OpenStudy (anonymous):

Calculate the radius of convergence and the interval of convergence (with examination of the endpoints) for the following power series: \sum_{n=0}^\infty \frac{n-1}{n+1} x^{n}

OpenStudy (anonymous):

Need help guys

OpenStudy (amistre64):

the radius and interval are 2 sides of the same coin

OpenStudy (amistre64):

find the limit as n goes to 0 of the ration an/a(n-1)

OpenStudy (anonymous):

sorry if i do something wrong i am quite new here, i need some solution, and sorry for my english

OpenStudy (amistre64):

the solution will come when we work thru the steps; this is a hands on study group

OpenStudy (amistre64):

ill guide you along the way :)

OpenStudy (anonymous):

ok thanks but i am doing it since 3 hours and i am without any solution and idea :(

OpenStudy (amistre64):

yeah, lets start at the beginning

OpenStudy (amistre64):

start by setting up the limit an/a(n-1)

OpenStudy (amistre64):

\[\lim_{n\to\ inf} \frac{(n-1)\ x^{n}}{n+1}*\frac{n-1+1}{(n-1-1)\ x^{n-1}} \] and simplify

OpenStudy (anonymous):

cool

OpenStudy (amistre64):

\[\lim_{n\to\ inf} \frac{(n-1)\ x}{n+1}*\frac{n}{(n-2)}\] pull out all the stuff that doesnt have an n attached to it \[|x|\ \lim_{n\to\ inf} \frac{(n-1)}{n+1}*\frac{n}{(n-2)}\] \[|x|\ \lim_{n\to\ inf} \frac{n^2...}{n^2...}=|x|\]

OpenStudy (amistre64):

so, lets set this up into your convergence: |x| < 1 ; there is no need to simplify since its as simple as we can get, so your radius = 1 and the interval is: -1 < x < 1 and it didnt even take me 1 hour to deal with

OpenStudy (amistre64):

but the only way to get good at it is to get your hands dirty in here

OpenStudy (anonymous):

:) wow, i think you are real mathematicer

OpenStudy (anonymous):

is it end solution like that?

OpenStudy (anonymous):

i write it all and try to understand each step

OpenStudy (amistre64):

i figure when we get to the end, its best to stop :) the wolf even agrees http://www.wolframalpha.com/input/?i=sum+%28n-1%29%2F%28n%2B1%29+x%5E%28n%29

OpenStudy (amistre64):

what that means is that the series matches the given function perfectly, it converges with the function, between -1 and 1

OpenStudy (anonymous):

ok, is wolfram giving the hole solution? by me it doesnt do it

OpenStudy (amistre64):

its not giving a step by step, but its affirming the outcome we reached. |x| < 1

OpenStudy (anonymous):

aah ok by me its doesnt do it, i am very thankful too you, its my first 10 min here and its realy cool website, and people here thank you very much

OpenStudy (amistre64):

youre welcome, and good luck. Word of advice, there are people who want to just hand you an answer, dont trust them. Always have them verify their results :)

OpenStudy (anonymous):

ok i will follow your advice many thanks, now i will try to understand your solution can i ask you question later if i stuck somewhere?

OpenStudy (amistre64):

i might not be around, but there should be someone who can give you a suitable, intelligent response. just post it up on the left like you did this one. You can only ask one question at a time to limit those who are trying to use this site to cheat with. So go ahead and press the close option to ask a new question.

OpenStudy (anonymous):

ok tthank you soo much i am very grateful.

OpenStudy (anonymous):

Let |x|<1 \[ \sum _{n=0}^{\infty } \frac{(n-1) x^n}{n+1}=\sum _{n=0}^{\infty } x^n-\sum _{n=0}^{\infty } \frac{2 x^n}{n+1}=\\ \frac 1 {1-x} -\frac 2 x \sum _{n=0}^{\infty } \frac{ x^{n+1}}{n+1}=\\ \frac 1 {1-x} -\frac 2 x (-\ln(1-x))=\\ \frac 1 {1-x}+\frac 2 x\ln(1-x) \\ \]

OpenStudy (anonymous):

I only found the sum of your series as the difference of two very well known series. This way, I was able to write the sum as a closed form function. You do not need to do that for your problem. It is nice to know the sum if it is feasible.

OpenStudy (anonymous):

ok thank you very much can i use your solution or above solution from amistre64 in my homework as a answer to this question? Or i need to additional calculation?

OpenStudy (anonymous):

you should do what @amistre64 suggested. My post is useful if you were asked to find the sum of that series.

OpenStudy (anonymous):

ok thank you very much